OpenStudy (anonymous):

The expression -3m - [2m + (5 - m)] + 7 was simplified as 2 - 4m. Without simplifying, explain how you can show that it has been simplified correctly.

jhonyy9 (jhonyy9):

so -3m-2m-(-m)=-5m+m=-4m and -(+5)+7=-5+7=2 so this mean that result -4m+2 ok ?

jhonyy9 (jhonyy9):

or 2-4m